# Privacy-Preserving Retrieval-Augmented Generation with Differential Privacy

### v2

@@ B.1 Main Results across Models and Datasets
B.2 Further Results on Trivia Dataset with OPT (1.3B)
Effects of Number of Ground Truth Relevant Documents.
Figure 7: Average match accuracy for questions with different numbers of ground truth relevant documents across baseline algorithms and our algorithms with OPT (1.3B) on Trivia dataset. For VoteRAG, DPVoteRAG, and DPSparseVoteRAG, we report the means and standard deviations over three runs.
Figure 7 shows the performances for different numbers of ground truth relevant documents. We see questions with more relevant documents tend to be answered correctly across algorithms, even with Non-RAG. We also observe that the benefit of having more voters in DPSparseVoteRAG (larger $m$) is larger for small $\epsilon_{\mathrm{token}}$. We don’t see significant difference between $m=30$ and $50$ when $\epsilon_{\mathrm{token}}=5$.
Number of Generated Tokens.
Figure 8: Average numbers of generated tokens for each $\epsilon_{\mathrm{token}}$ and $\epsilon_{\mathrm{total}}$ with OPT (1.3B) on Trivia dataset. We fix $m=50$. We report the means and standard deviations over three runs.
Figure 8 shows the numbers of tokens generated by DPVoteRAG and DPSparseVoteRAG. As we expect by the design of DPSparseVoteRAG, we see DPSparseVoteRAG generates much more tokens than DPVoteRAG. This implies the effectiveness of the sparse vector technique in DPSparseVoteRAG to smartly spend privacy budget enabling long enough token sequences.

### v3

@@ VI Conclusion and Future Work
TABLE XIII: Average match accuracy comparison of our algorithms on ChatDoctor dataset with OPT (1.3B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.82$ $\mathbf{0.83}$ $\mathbf{0.83}$ $0.82$ $0.83$ $0.83$ $0.82$ $\mathbf{0.83}$ $\mathbf{0.83}$ $0.82$ $\mathbf{0.83}$ $\mathbf{0.83}$
$\epsilon_{\mathrm{token}}=2$ $0.82$ $0.82$ $0.81$ $0.83$ $0.83$ $0.83$ $\mathbf{0.83}$ $\mathbf{0.83}$ $\mathbf{0.83}$ $\mathbf{0.83}$ $\mathbf{0.83}$ $\mathbf{0.83}$
$\epsilon_{\mathrm{token}}=5$ $0.82$ $0.82$ $0.82$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.84}$ $0.82$ $0.82$ $0.82$ $\mathbf{0.83}$ $\mathbf{0.83}$ $\mathbf{0.83}$
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.79$ $0.80$ $0.81$ $0.79$ $0.80$ $0.81$ $0.79$ $0.80$ $0.81$ $0.79$ $0.80$ $0.81$
$\epsilon_{\mathrm{token}}=2$ $0.83$ $\mathbf{0.84}$ $\mathbf{0.84}$ $0.82$ $\mathbf{0.84}$ $\mathbf{0.84}$ $0.82$ $0.84$ $\mathbf{0.85}$ $0.82$ $0.84$ $\mathbf{0.85}$
$\epsilon_{\mathrm{token}}=5$ $0.13$ $0.13$ $0.13$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.84}$ $0.84$ $0.84$ $0.84$ $\mathbf{0.85}$ $\mathbf{0.85}$ $\mathbf{0.85}$
TABLE XIV: Average match accuracy comparison of our algorithms on ChatDoctor dataset with Pythia (1.4B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.82$ $0.84$ $\bm{0.85}$ $0.82$ $0.84$ $\bm{0.85}$ $0.82$ $0.84$ $\bm{0.85}$ $0.82$ $0.84$ $\bm{0.85}$
$\epsilon_{\mathrm{token}}=2$ $0.82$ $0.82$ $0.82$ $0.84$ $\bm{0.85}$ $\bm{0.85}$ $0.84$ $\bm{0.85}$ $\bm{0.85}$ $0.84$ $\bm{0.85}$ $\bm{0.85}$
$\epsilon_{\mathrm{token}}=5$ $0.84$ $0.84$ $0.84$ $0.81$ $0.82$ $0.81$ $0.84$ $0.84$ $0.84$ $\bm{0.85}$ $\bm{0.85}$ $\bm{0.85}$
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.79$ $0.80$ $0.81$ $0.79$ $0.80$ $0.81$ $0.79$ $0.80$ $0.81$ $0.79$ $0.80$ $0.81$
$\epsilon_{\mathrm{token}}=2$ $0.82$ $0.82$ $\bm{0.83}$ $0.83$ $0.84$ $\bm{0.85}$ $0.83$ $\bm{0.85}$ $\bm{0.85}$ $0.83$ $\bm{0.85}$ $\bm{0.85}$
$\epsilon_{\mathrm{token}}=5$ $0.82$ $0.82$ $0.82$ $0.83$ $0.83$ $0.83$ $\bm{0.85}$ $\bm{0.85}$ $\bm{0.85}$ $\bm{0.85}$ $\bm{0.85}$ $\bm{0.85}$
TABLE XV: Average match accuracy comparison of our algorithms on ChatDoctor dataset with Llama 3.1 (8B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.83$ $\mathbf{0.84}$ $\mathbf{0.84}$ $0.83$ $\mathbf{0.84}$ $\mathbf{0.84}$ $0.83$ $\mathbf{0.84}$ $\mathbf{0.84}$ $0.83$ $0.84$ $0.84$
$\epsilon_{\mathrm{token}}=2$ $0.83$ $0.83$ $0.83$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.84}$ $0.84$ $0.84$ $0.84$
$\epsilon_{\mathrm{token}}=5$ $0.82$ $0.82$ $0.82$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.85}$ $0.84$ $0.84$
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.82$ $0.83$ $0.82$ $0.82$ $0.83$ $0.82$ $0.82$ $0.83$ $0.82$ $0.82$ $0.83$ $0.82$
$\epsilon_{\mathrm{token}}=2$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.84}$ $\mathbf{0.85}$ $\mathbf{0.85}$ $\mathbf{0.85}$ $0.85$ $0.85$ $0.85$ $0.85$ $0.85$ $0.85$
$\epsilon_{\mathrm{token}}=5$ $0.46$ $0.45$ $0.35$ $0.84$ $\mathbf{0.85}$ $0.84$ $0.85$ $\mathbf{0.86}$ $0.85$ $\mathbf{0.86}$ $\mathbf{0.86}$ $\mathbf{0.86}$